Here's a problem with equations and tables.(TABLE): Months(x)0 2 3 4
Money in Bank(y)25 33 41 49 57
The question is: The table shows the amount of money Betty has in her bank account. Which equation on the left matches the table on the right?
Options to choose from: A y=8x B y=1x C y=8x+55 D y=1x+55
- -
8 8

Answer by stanbon(75887)

Here's a problem with equations and tables.(TABLE): Months(x)0 2 3 4
Money in Bank(y)25 33 41 49 57
The question is: The table shows the amount of money Betty has in her bank account. Which equation on the left matches the table on the right?
Options to choose from: A y=8x B y=(1/8)x C y=8x+55 D y=(1/8)x+55
----
Note: You have 4 x-values and 5 y-values
None of those answers match the data.
---
The data seems to imply y = 8x+25
=======================================
Cheers,
Stan H.